Factor completely. 10x2 + 10x – 200

Found 2 solutions by checkley71, monica17:
Answer by checkley71(8403)

(5X-20)(2X+10)=0 OR 5X-20=0 OR 54X=20 OR X=4 & 2X+10=0 OR 2X=-10 OR X=-5

Answer by monica17(5)

Factor completely: 10x^2+10x-200
The first thing you ought to do is:
factor out a common factor
10(x^2 +x -20)
= 10(x+5) (x-4)
Let me know if you need further explanation.